The mass percent composition of a particular compound was found to be 26.95% sulfur, 14.43% oxygen, and 59.61% chlorine. The molar mass of the compound was determined to be 118.96 g·mol−1118.96 g·mol−1.

Determine the molecular formula for this compound.

S2OCl6S2OCl6

S2O2Cl4S2O2Cl4

SOClSOCl

S2O2Cl2S2O2Cl2

SOCl2
Answer #1

Consider mass of compound is 100 g

No. of moles of S = (26.95 g) / (32 g/mol) = 0.842 mol

No. of moles of O = (14.43 g) / (16 g/mol) = 0.9 mol

No. of moles of Cl = (59.61 g) / (35.45 g/mol) = 1.681 mol

Simplest mol ratio is:

S = 0.842 mol / 0.842 mol = 1

O = 0.9 mol /0.842 mol = 1 (approx)

Cl = (1.681 mol) / (0.842 mol) = 2

Empirical formula is: SOCl2

Molar mass of emipirical formula is: 118.96 g/mol

Here, empirical mass is equals to molar mass

So, Compound molecular formula is: SOCl2